Determine the quadratic equation whose solutions are: 3 and -2

Answer by Maths68(1474)

Solution set = 3, -2
x= 3, or -2
x=3 or x=-2
x-3=0 or x+2=0
Since either of above is equal to zero, therefore multiplicatoin of the both will equal to zero
(x-3)(x+2)=0
x^2+2x-3x-6=0
x^2-x-6=0
Required Quadratic Equation.